OpenStudy (anonymous):

Use the given values to evaluate all six trigonometric functions. cos (pi/2-x)=3/5, cosx=4/5

OpenStudy (cwrw238):

Cos (pi/2 - x) = sin x = 3/5 tan x = sinx / cos x = (3/5) / (4/5) sec x = 1 / cos x, csc x = 1 / sinx cot x = 1 / tan x